OpenStudy (anonymous):

Find the value of x so that the line passing through (-2, 7) and (x, 2) has a slope of

OpenStudy (anonymous):

one third

geerky42 (geerky42):

Use slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) We have \(x_1 =-2,~~x_2 =x,~~y_1 =7,~~y_2 =2\), right? So you have \(\dfrac{1}{3} = \dfrac{2-7}{x-(-2)}\) Try do cross-multiply then solve for x.

OpenStudy (anonymous):

13?

geerky42 (geerky42):

I got -17 ? \(3\cdot-5 = x+2\\~\\-15 = x+2\\~\\x = -17\)

OpenStudy (anonymous):

I did -0.33333333333333(x-(-2))=2-7 -0.33333333333333x-0.66666666666666=-5 add 0.66666666666666 to both sides -0.33333333333333x=-4.33333333333334 divide -0.33333333333333 from both sides x=13

OpenStudy (anonymous):

sorry I used decimals instead of fractions so there are a lot of numbers :/

geerky42 (geerky42):

Wait, slope is \(-\dfrac{1}{3}\), not \(\dfrac{1}{3}\)? You just said one third.

OpenStudy (anonymous):

ya sorry it was negative I think :/